### v3

# Wenfan Ou and Sheng Bi

[…]
Wenfan Ou and Sheng Bi
[…]
Sequential Decision-Making under Uncertainty: A Robust MDPs review
[…]
Fueled by both advances in robust optimization theory and applications of reinforcement learning, robust Markov Decision Processes (RMDPs) have gained increasing attention, due to their powerful capability for sequential decision-making under uncertainty. This review provides an in-depth overview of the evolution and advances in RMDPs formulations, particularly in ambiguity modeling, and classifies these methods for representing uncertainty into three principal approaches: parametric, moment-based, and discrepancy-based, elaborating the trade-offs among the alternative representations. Meanwhile, the review delves into the rectangular assumptions, which guarantee the tractability of RMDPs yet are noted for their conservatism. The review summarizes three popular rectangular conditions and develops a new proof to attest to the NP-hardness of non-rectangular RMDPs. Out of the traditional RMDPs scope, recent efforts without conventional rectangular assumptions and new fashions within the RMDPs community are also reviewed. These studies foster the development of more flexible and practical modeling frameworks and enhance the adaptability and performance of RMDPs.
[…]
robust Markov Decision Processes, ambiguity sets modeling, reinforcement learning, rectangularity, new fashions
1 Introduction
Markov decision processes (MDPs) originate from the seminal works of BELLMAN (1957) and Howard (1960) to model sequential decision-making problems under uncertainty. As powerful mathematical tools, MDPs are widely applied in many areas, such as economics, transportation, inventory control, health care, and manufacture (Parkes & Singh 2003, McKenna et al. 2020, Song & Zipkin 1993, Katehakis & Puranam 2012, Fan et al. 2022, Che et al. 2023). Recently, reinforcement learning (RL), which establishes its theoretical foundations on MDPs (Sutton & Barto 2018), has attracted increasing attention and achieved notable success in practical applications. RL stands out for its potential intelligence on decision-making and active learning capabilities, especially when incorporating deep learning (Matsuo et al. 2022, Kosanoglu et al. 2022, Bushaj et al. 2023, Rolf et al. 2023, Zuccotto et al. 2024). Undeniably, the advances in RL pave the way for wider applications of MDPs. However, volatile and dynamic environments always cause poor performance of MDPs for sequential decision-making. Naive MDPs presuppose a complete grasp of the underlying dynamics, an assumption that often falls short in the real world. With deviated estimation or belief affected by uncertainty, decision-makers should anticipate disappointment on average, owing to the optimization-based selection process. This is the so-called the optimizer’s curse (Smith & Winkler 2006). Even worse, this skewed belief in sequential decision-making will persist throughout the entire planning period, persistently leading to worse outcomes compared to single-stage decision-making problems.
Motivated by advances in robust optimization theory and practical performance (Scarf et al. 1957, Ben-Tal & Nemirovski 2002, Bertsimas & Sim 2004, Goerigk & Schöbel 2016), robust Markov decision processes (RMDPs) have been investigated for decades. Unlike conventional MDPs, RMDPs are specifically designed to deal with situations where only partial information about the true underlying parameters or probability distributions is available, rather than complete knowledge.
The decision-maker can construct a collection of possible parameter values or distributions, called an uncertainty set or ambiguity set, based on the belief that the partial information is credible. Subsequently, the decision-maker can determine the optimal solution under the worst-case scenario over the uncertainty or ambiguity set.
The exploration of RMDPs can be traced back to the 1970s (Satia & Lave Jr 1973) when they were known as MDPs with imprecisely known parameters (MDPIPs) (White III & El-Deib 1986, White III & Eldeib 1994, Givan et al. 2000). Most MDPIPs assume the transition probabilities are bounded by linear programs, appealing to linear programming at each iteration. While defining the uncertainty on the transition matrix by a polytope is natural and convenient, the approach brings some drawbacks. Polytopic models not only pose computational challenges but also lead to excessively conservative policies due to the possibly terrible representations of statistical uncertainty. To address this issue, seminar papers by Iyengar (2005) and Nilim & El Ghaoui (2005) independently propose a similar formulation that leverages strong duality proofs and rectangularity assumption to extend value iteration for MDPs to robust value iteration. This formulation is also referred to as robust DP, and serves as a modern incarnation of RMDPs thereafter. Both two papers introduce two imperative concepts. Firstly, they introduce ambiguity sets, especially $\phi$-divergences, norm, and likelihood region ambiguity sets, which outperform polytopic models. Secondly, these works coincidentally emphasize that the rectangularity assumption is essential for the tractability of RMDPs, albeit without providing concrete proof.
To fill the theoretical gap, Bagnell et al. (2001) and Wiesemann et al. (2013) provide the NP-hard proofs for general (non-rectangular) RMDPs in different ways, respectively. Moreover, with the conservatism of robust paradigm proved widely (Thiele 2010, Delage & Mannor 2010) and innovations of distributionally robust optimization (DRO) (Popescu 2007, Delage & Ye 2010, Goh & Sim 2010), Xu & Mannor (2012) first propose the framework of distributionally robust MDPs (DRMDPs). Compared to traditional RMDPs, the new framework is more convenient to incorporate statistical information to mitigate conservatism. Xu & Mannor (2012) and Yu & Xu (2015) show that DRMDPs can be reduced to expected standard RMDPs under some assumptions, thereby retaining the tractability. These research findings have laid a solid theoretical foundation for RMDPs, and enable the thriving development of the RMDP community.
Beneficial from substantial results across diverse fields and interests in data-driven practices, the literature on RMDPs has been significantly enriched in recent years. Researchers have explored RMDP modeling from various aspects, including the exploration of diverse ambiguity sets, the development of effective solving and approximation algorithms, as well as the design of robust algorithms for RL. Furthermore, RMDPs have been combined with other techniques from various perspectives and applied in many areas. These techniques include online learning (Croonenborghs et al. 2007, Badrinath & Kalathil 2021, Cowan et al. 2018), safe learning (Hans et al. 2008, Polo & Rebollo 2011, Wachi & Sui 2020), dynamic risk measure (Bäuerle & Rieder 2014, Chow et al. 2015, Yu & Shen 2022, Rockafellar & Uryasev 2000), and regularization (Farahmand 2011, Zhang et al. 2018), etc. Despite the surge in the literature on RMDPs in recent years, a systematic review of RMDPs formulation remains scarce, with only a handful of tutorials or general reviews available (Mannor & Xu 2019, Keith & Ahner 2021, Badings et al. 2023). We believe that the absence of a specialized review hinders the development of RMDP theory and applications. To bridge this gap, we provide an up-to-date review of the RMDPs/DRMDPs formulations with uncertain transitions. This is a common class of RMDPs problems that have been widely considered in research and applications within the management science and operations research community. By surveying impressive results and studies in this domain, we aim to establish a solid foundation for future research and practical implementations in RMDPs.
The remainder of the paper is organized as follows. In Section 2, we briefly introduce the preliminary of RMDPs. In particular, we summarize the definitions of rectangularity and provide a new proof for NP-hard of non-rectangular RMDPs. In Section 3, 4 and 5, we respectively introduce three popular types of RMDPs: parametric RMDPs, moment-based RMDPs, and discrepancy-based RMDPs. In these sections, we focus on generic formulations and key theoretical advancements. Although rectangularity maintains the tractability of RMDPs, it often doesn’t hold in practice. In Section 6, we investigate methodologies for modeling coupled uncertainty that violates rectangularity assumptions. Finally, we review some novel works in Section 7 that have sparked our interest beyond the traditional mini-max framework and conclude in Section 8.
2 Preliminaries
We denote by $[T]=\{1,2,...,T\}$ the set of positive running indices up to $T$, and use $\bigotimes$ to represent the Cartesian product. Let $\mathfrak{M}(\Xi)$ be a set of probability measures on measurable space $(\Xi,\mathcal{F})$, where $\Xi:=\bigotimes_{t\in[T]}\Xi_{t}$ is finite, and $\Xi_{t}$ is the sample space at stage $t$. Define $\left\{\emptyset,\Xi\right\}=\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\cdots\subseteq\mathcal{F}_{T}=\mathcal{F}$. Let $\mathcal{P}\subseteq\mathfrak{M}(\Xi)$ be the set of probability measures we are interested in, and $P\in\mathcal{P}$ be a probability measure on $(\Xi,\mathcal{F})$. Without loss of generality, we can appropriately derive the corresponding probabilities $\mathbf{p}$ or a (discrete) distribution $\mu$ with respect to the probability measure $P$. Let $\mathcal{P}_{t}(\xi)=\left\{P(\cdot|\mathcal{F}_{t})(\xi)\ |\ P\in\mathcal{P}\right\}$. $\Delta^{d}=\left\{\mathbf{p}\ |\ p_{1}+\cdots+p_{d}=1,p_{i}\geq 0\ \forall i\in[d]\right\}\in\mathbb{R}^{d}$ is the probability simplex. $\bm{1}(\cdot)$ is an indicator function. We use superscript 00 to denote true underlying value, e.g., $\mathbf{p}^{0}$, and $\hat{\cdot}$ denotes the estimated or nominal one, e.g., $\hat{\mathbf{p}}$.
[…]
$\max_{\pi}\mathbb{E}_{\mathbf{p}}^{\pi}\left[{\textstyle\sum\limits_{t=1}^{T}\gamma^{t-1}R(s_{t},a_{t},s_{t+1})+\gamma^{T}R(s_{T+1})}\right]$ (1)
where $0<\gamma\leq 1$ is the discounted factor. Consequently, the (randomized) policy $\pi:\mathcal{S}\rightarrow\Delta^{|\mathcal{A}|}$ specifies a decision-making strategy in which the decision-maker chooses actions adaptively based on the current state, i.e., takes action $a\in\mathcal{A}$ with probability $\pi(s,a)$ at state $s\in\mathcal{S}$. If $\pi:\mathcal{S}\rightarrow\mathcal{A}$, we call it a deterministic policy.
Let $V_{t}(\cdot):\mathcal{S}\rightarrow\mathbb{R}$ denote the value-to-go function at stage $t$, solving (1) is equivalent to solving the following equations recursively:
$V_{t}(s_{t})=\max_{\pi}\mathbb{E}_{\mathbf{p}}^{\pi}\left[R(s_{t},a_{t},s_{t+1})+\gamma V_{t+1}(s_{t+1})\right]$ (2)
where $V_{T+1}$ is the terminal value which takes the value of 0 or $-R(s_{T+1})$ conventionally.
The value-to-go function (2) can be solved efficiently by dynamic programming techniques (finite horizons) and value/policy iteration methods (infinite horizons) (Puterman 2014), although there exists the so-called “curse of dimensionality” (Bellman 1966).
Note that with the growing research on MDPs, there have been diverse representations of MDPs proposed, such as assuming known initial state distributions or continuous state spaces. However, the essential elements of MDPs outlined above capture the core components and remain widely applicable across this review, unless otherwise specified. Future extensions can build on this general MDP formulation by incorporating additional details and specifications tailored to particular problems.
Remark 2.1
The transition probability matrix $\mathbf{p}$ defined above is associated with the transition tuple $(s,a,s^{\prime})$, but most literature assumes exogenous randomness. For example, in inventory management problems, the demand is often not affected by inventory policies. Therefore, $\mathbf{p}$ may be independent of actions and can be represented by a distribution directly.
2.1 (Distributionally) Robust Markov Decisions Processes
As mentioned in Section 1, we focus on MDPs with uncertain transition probability. To ensure generality, we claim that the set $\mathcal{P}$ consists of probability measures on probability space. However, we sometimes abuse the notation $\mathcal{P}$ to represent the set of corresponding transition probabilities $\mathbf{p}$ or distribution $\mu$ to align with literature. Formally, we consider RMDPs and DRMDPs, defined as follows:
Definition 2.2 (Robust MDPs)
A RMDP is defined as a tuple $<T,\mathcal{S},\mathcal{A},\mathcal{P},\gamma>$, where the transition probability matrix $\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}$ is unknown but belongs to a known collection $\mathcal{P}$ of parameters, called uncertainty set. Exactly, RMDPs can be expressed as
$\max_{\pi}\min_{\mathbf{p}\in\mathcal{P}}\mathbb{E}^{\pi}_{\mathbf{p}}\left[{\textstyle\sum\limits_{t=1}^{T}\gamma^{t-1}R(s_{t},a_{t},s_{t+1})+\gamma^{T}R(s_{T+1})}\right].$ (3)
Definition 2.3 (Distributionally robust MDPs)
A DRMDP can be represented by a tuple $<T,\mathcal{S},\mathcal{A},\mathcal{P},\gamma>$, where the transition probability matrix $\mathbf{p}$ obeys to a distribution $\mu$, denoted by $\mathbf{p}\sim\mu$. $\mu$ is unknown but belongs to a known distribution family $\mathcal{P}$ of probability distributions, called the ambiguity set. Exactly, DRMDPs can be expressed as
$\max_{\pi}\min_{\mu\in\mathcal{P}}\mathbb{E}^{\pi}_{\mathbf{p}\sim\mu}\left[{\textstyle\sum\limits_{t=1}^{T}\gamma^{t-1}R(s_{t},a_{t},s_{t+1})+\gamma^{T}R(s_{T+1})}\right].$ (4)
While RMDPs view the transition probabilities as uncertain parameters, DRMDPs model them as random variables or use the distribution of exogenous randomness to represent dynamics. This difference in definitions is noteworthy, bringing distinct formulation styles and enabling distinct ambiguity set modeling. For example, moment-based ambiguity sets are more prevalent in DRMDPs compared to RMDPs.
Despite the differences between RMDPs and DRMDPs, they exhibit similar formulations within the scope of this article. The high-level modeling approaches exhibit substantial parallels, and both formulations can be converted to each other with appropriate modeling and practical undergrounds, especially in a discrepancy-based framework (Section 5). Thus, this article mainly investigates the literature based on modeling frameworks, rather than specifically differentiating RMDPs and DRMDPs. Unless otherwise noted, RMDPs and DRMDPs will both be referred to as RMDPs in the following for simplicity.
2.2 Rectangularity
While the definition of RMDPs is intuitive, the tractability of MDPs cannot be straightforwardly extended to RMDPs, as it heavily relies on the rectangularity assumption. Before diving into the importance of rectangularity for RMDPs, we first deliver the definition of this strict assumption. However, various communities have studied several closely related problems at different times, and a complete consensus on a common rigorous definition has not yet been reached (Xin & Goldberg 2021). We present several popular definitions to convey the core idea of rectangularity.
Definition 2.4 (Rectangularity)
Lending the mathematical representations, the definition of rectangularity can be classified into three types:
(1) Static form (Nilim & El Ghaoui 2005, Le Tallec 2007, Wiesemann et al. 2013). In this case, $s$-rectangular ambiguity set and $(s,a)$-rectangular ambiguity set are time-invariant. They can be represented respectively as follows:
$\mathcal{P}^{\mathcal{S}}$ =$=\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\ |\ \mathbf{p}=\bigotimes_{s\in\mathcal{S}}\mathbf{p}_{s},\mathbf{p}_{s}\in\mathcal{P}_{s},\mathcal{P}_{s}\subseteq\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\right\},$
$\mathcal{P}^{\mathcal{SA}}$ =$=\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\ |\ \mathbf{p}=\bigotimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbf{p}_{sa},\mathbf{p}_{sa}\in\mathcal{P}_{sa},\mathcal{P}_{sa}\subseteq\Delta^{|\mathcal{S}|}\right\}.$
(2) Dynamic form (Iyengar 2005, Epstein & Schneider 2003, Iancu et al. 2015). Let $P^{+1}(\cdot|\mathcal{F}_{t})(\xi)$ be the restriction of $P(\cdot|\mathcal{F}_{t})(\xi)$ to $\mathcal{F}_{t}+1$. In this case, the ambiguity set is rectangular if it satisfies:
$\mathcal{P}_{t}(\xi)=\left\{\int P(\cdot|\mathcal{F}_{t+1})(\xi^{\prime})\mathrm{d}Q(\xi^{\prime})\ |\ P\in\mathcal{P},\forall\xi^{\prime}\in\Xi_{t+1},Q\in\mathcal{P}_{t}^{+1}(\xi)\right\},$
where $\mathcal{P}_{t}^{+1}(\xi)=\left\{P^{+1}(\cdot|\mathcal{F}_{t})(\xi)\ |\ P\in\mathcal{P}\right\}$.
(3) Nested form (Shapiro 2016). Let $f(\xi_{[T]})=\sum_{t=1}^{T}\gamma^{t-1}R(s_{t},a_{t},s_{t+1})+\gamma^{T}R(s_{T+1})$ for simplicity. In this case, the rectangular assumption is satisfied if the following equation holds,
$\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left[f(\xi_{[T]})\right]=\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left[\sup_{P\in\mathcal{P}}\mathbb{E}_{P|\xi_{1}}\left[\cdots\sup_{P\in\mathcal{P}}\mathbb{E}_{P|\xi_{[T-1]}}[f(\xi_{[T]})]\cdots\right]\right].$
The first form is termed “static” because it lacks a subscript $t$. The static form of rectangularity aims to connect the state-wise independence with rectangularity, and a more concise representation is given by $\mathcal{P}^{\mathcal{S}}=\bigotimes_{s\in\mathcal{S}}\mathcal{P}_{s}$. It is noteworthy that the static form is often the most popular representation of rectangularity in RMDPs literature. In the risk measure community, the dynamic form of the rectangular structure is derived using one-step conditional measures, connecting rectangularity with time-(in)consistency. Recently, Shapiro (2016) proposed a nested-form condition that ensures the rectangularity assumption, directly targeting the decomposability property of the static formulation to establish the equivalence with the dynamic formulation.
While there are interesting discussions on the relationship between rectangularity and time-(in)consistency, they are beyond the scope of this review. We refer interested readers to Xin & Goldberg (2021) and the references therein for details.
Remark 2.5
It is important to note that these definitions of rectangularity are not simply semantically distinct, their underlying meanings differ fundamentally as well. Though definitions differ, stage-wise independence stands as a common condition that can satisfy rectangularity as characterized in any formulation.
In the context of game theory, $(s,a)$-rectangular ambiguity sets and $s$-rectangular ambiguity sets differ in their interpretation of nature’s knowledge and commitment (Le Tallec 2007). $(s,a)$-rectangular ambiguity sets essentially assume that nature can observe the decision-maker’s actions before choosing the worst plausible realization of the transition probabilities. In other words, nature can react to the decision-maker’s actions and choose the worst-case scenario among all possible realizations of the transition probabilities that are consistent with the observed actions. This interpretation is often referred to as a reacting or adaptive nature. While $s$-rectangular ambiguity sets assume that a weaker nature must commit to a realization of the transition probabilities before observing the decision-maker’s actions. Hence, nature must choose a realization of the transition probabilities without observing the decision-maker’s actions, and this realization must be worst-case over all possible realizations that are consistent with the ambiguity sets. While different rectangular assumptions adapt to different situations, $s$-rectangular ambiguity sets derive less conservative solutions in general.
Rectangularity assumption has been investigated widely about the tractability of RMDPs, and Bagnell et al. (2001) and Wiesemann et al. (2013) prove that solving a general (non-rectangular) RMDP is NP-hard. We provide a new proof that is similar to Bagnell et al. (2001) by constructing a non-rectangular RMDP as a conjunctive normal form (CNF) formula yet more accessible. In addition, we also show how to add the rectangularity condition such that the RMDPs can be solved in polynomial time.
Proposition 2.6 (NP-hard for non-rectangularity)
Solving the non-rectangular RMDPs is NP-hard.
Proof 2.7
The CNF formula provided is used to create a non-rectangular and simplified RMDP $<\mathcal{S},\mathcal{A},\mathbf{p}>$ without discount and reward randomness. Each clause $i$ is defined as state $s_{i}$, while variable $x_{j}$ in each clause $i$ is defined as a reachable state $s_{j}$. For each $x_{j}$ in clause $i$, “True” is assigned with probability $p_{ij}$, and “False” is assigned with probability $1-p_{ij}$. Note that $p_{ij}$ is binary for a deterministic policy $\pi$. As a deterministic policy $\pi$ is selected, the worst-case scenario must be deterministic. Therefore, if “True” is assigned to variable $x_{j}$ in clause $i$, the state will move to $s_{j}$ from $s_{i}$ after taking action.
If a policy $\pi$ developed above can solve a non-rectangular robust MDP in polynomial time, it can be used to solve the CNF-SAT problem in polynomial time. However, it is well-known that solving the CNF-SAT problem is NP-hard. Therefore, we prove that solving a non-rectangular RMDP is NP-hard.
It is noteworthy that a non-rectangular RMDP may be transformed into a rectangular RMDP by introducing an additional constraint. In the general rectangular RMDPs formulation, each state is visited only once, and each visit is only to one state, making each state distinct from the others despite their possibly identical “sensation”. In the CNF setting, the variables in each clause differ from those in the other clauses, allowing for decomposition and the creation of a policy that assigns values to the variables in each clause individually. This can be solved in polynomial time, highlighting the significance of the rectangular assumption in RMDPs research.
Finally, we briefly present several novel rectangular concepts to close this section. Mannor et al. (2016) propose the concept of $k$-rectangularity, which restricts the cardinality of deviations to no more than $k$ to maintain tractability. Goh et al. (2018) and Goyal & Grand-Clément (2023) decompose transition probabilities into linear combinations of $r$ factors. Essentially, they require independence of factors, instead of probabilities, and term the new assumption as $r$-rectangularity. More details about these novel rectangular concepts will be specified in Section 6. While beyond traditional rectangular assumptions, these concepts build upon the rectangularity class, which remains fundamental to RMDPs.
3 Parametric RMDPs
Parametric RMDPs are among the earliest studied RMDPs (Satia & Lave Jr 1973), which mainly assume that known parameters control transition probabilities. By predetermining parameters supports or distribution families, parametric RMDPs construct uncertainty sets in a parametric fashion. The study on parametric RMDPs has undergone significant changes, even though they fall into the same category in this paper. The development of RMDPs has seen major shifts, leading to the exploration of new approaches and techniques in parametric RMDPs.
As a pioneering work, Satia & Lave Jr (1973) propose an unprecedented formulation, known as MDPIPs, where ambiguity sets comprise individually bounded transition probabilities with given upper/lower bounds. The uncertainty sets $\mathcal{P}$ are defined by:
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\left|\begin{array}[]{l}p^{L}_{sas^{\prime}}\leq p_{sas^{\prime}}\leq p^{U}_{sas^{\prime}},\\ \textstyle\sum_{s^{\prime}}p_{sas^{\prime}}=1,\\ p_{sas^{\prime}}\geq 0,~{}\forall s,s^{\prime}\in\mathcal{S},a\in\mathcal{A}\end{array}\right\}\right.,\$ (5)
where $p^{L}_{sas^{\prime}}$ and $p^{U}_{sas^{\prime}}$ are the predetermined lower bound and upper bound, respectively. The uncertainty set $\mathcal{P}$ in (5) builds a polyhedron. Consequently, the authors develop a policy iteration algorithm that leverages linear programming techniques in each iteration to address the challenges posed by robust formulations effectively. They prove the proposed algorithm can find an $\epsilon$-optimal policy in a finite number of iterations.
In a continuous line, Givan et al. (2000) investigate a special subclass of MDPIPs, termed as bounded-parameter MDPs (BMDPs). A BMDP $M_{\updownarrow}=(T,\mathcal{S},\mathcal{A},R_{\updownarrow},\mathbf{p}_{\updownarrow},\gamma)$ is similar with the general MDP, but $\mathbf{p}_{\updownarrow}$ and $R_{\updownarrow}$ are closed real intervals instead of real values. Compared to previous work under the general setting, the authors consider a particular MDPs family as the order-maximizing MDPs. By utilizing the special structure of this family, they show that BMDPs bound the optimal value function and directly derive the optimal solution without requiring expensive optimization techniques.
More recently, Delimpaltadakis et al. (2023) also leverage the threshold structure to expand the scope of BMDPs into continuous action spaces. Distinct from simple real-valued intervals, the authors introduce two functions $f^{L},f^{U}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ to represent the lower and upper bounds on the transition probability with respect to transition $(s,a,s^{\prime})$. Due to strong duality and the ordering threshold structure, they successfully convert the original max-min formulation into a maximization problem. Furthermore, they develop an efficient value iteration algorithm based on this reformulation when $\mathcal{A}$ is a polytope
Besides defining the range of transition probabilities directly, Osogami (2012) considers parametric RMDPs with $s$-rectangular uncertainty sets specified by the factor, $10<\alpha<$ 1, which determines the possibly maximal value of each transition probability based on the nominal values:
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\left|0\leq p_{sas^{\prime}}\leq\frac{1}{\alpha}\hat{p}_{sas^{\prime}},\forall s^{\prime}\in\mathcal{S}\text{ and }\textstyle\sum_{s^{\prime}\in\mathcal{S}^{\prime}}p_{sas^{\prime}}=1\right\}\right..$ (6)
Contrary to explicitly defining intervals for each probability, this approach just needs to determine a unique parameter $\alpha$ such that the approach can more conveniently construct ambiguity sets and achieve better generalization performance when the decision-maker has limited information about transitions. Additionally, $\alpha$ can be interpreted as the degree of robustness or risk-aware confidence level. In this paper, the author establishes the equivalence between RMDPs and risk-sensitive MDPs, where the risk confidence level is set to $1-\alpha$. This relationship provides further insights to understand and construct ambiguity sets.
Contrary to the polytopic uncertainty sets, Wiesemann et al. (2013) consider ambiguity sets $\mathcal{P}$ of the following type:
$\mathcal{P}\triangleq\left.\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\right|\exists\ \theta\in\Theta\text{, such that }\mathbf{p}_{sa}:=P^{\theta}(\cdot\mid s,a),~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}\right\},$ (7)
where $\Theta$ is a subset of $\mathbb{R}^{q}$ and $P^{\theta}(\cdot\mid s,a),~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ is an affine function from $\Theta$ to $[0,1]$ that satisfies $P^{\theta}(\cdot\mid s,a):=k_{sa}+K_{sa}\theta$ for some $k_{sa}\in\mathbb{R}^{|\mathcal{S}|}$ and $K_{sa}\in\mathbb{R}^{|\mathcal{S}|\times k}$. Note that $\theta$ controls the ambiguity sets essentially, which allows to condense all ambiguous information and parameters in the set $\Theta$, e.g.,
$\Theta\triangleq\left\{\theta\in\mathbb{R}^{q}\colon\theta^{\top}O_{l}\theta+o_{l}^{\top}\theta+\omega\geq 0,~{}\forall l=1,\ldots,L\right\},$ (8)
where $O_{l}$ is a $k\times k$ matrix, $O_{l}\preceq 0$, $\omega$ is the constant and $L$ is the number of constraints applied to $\theta$. The benefit of adhering to this construction as defined in (8) is that the set $\Theta$ can be derived as the finite intersection of closed half-spaces and ellipsoids, enabling efficient policy iteration by tailoring to second-order cone programs and semi-definite programs. They assume that $\Theta$ is bounded and has a nonempty interior, which means none of the parameters in $\Theta$ is fully explained by the others. With the $s$-rectangular assumption, the authors show that solving an RMDP is equivalent to finding a fixed point of the robust Bellman operator, and provide the time complexity of policy evaluation and policy improvement, respectively.
Black et al. (2022) investigate RMDPs with the ambiguity sets which only contain distribution from the same parametric family, focusing on identifying the worst-case parameters instead of the entire distribution. This approach aims to alleviate challenges with moment estimation in non-parametric RMDPs, particularly the issue of overly conservative solutions that can arise when parametric distributions are fitted well enough. For each $(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathbf{p}_{sa}$ is uniquely defined by the distribution of some exogenous random variable $\xi_{sa}\in\Xi_{sa}$. Let $f_{\xi_{sa}}$ be the probability mass function of $\xi_{sa}$ which is controlled by the parameters $\bm{\theta}_{sa}=(\theta_{sa1},\cdots,\theta_{sak})$. The authors assume that the next state $s_{t+1}$ is specified by a simple, known function $g$ of the exogenous random variable $\xi_{sa}$ as $s_{t+1}=g(\xi_{sa}|s,a)$. Hence, the transition probability $p_{sas^{\prime}}$ for each transition tuple $(s,a,s^{\prime})$ can be computed as
$p_{sas^{\prime}}=P^{\bm{\theta}_{sa}}(s^{\prime}|s,a)=P^{\bm{\theta}_{sa}}(g(\xi_{sa}|s,a)=s^{\prime})=\sum_{\xi\in\Xi_{sa}(s^{\prime})}f_{\xi_{sa}}(\xi|\bm{\theta}_{sa}),$
where $\Xi_{sa}(s^{\prime})$ is the support of $\xi$ with $(s,a,s^{\prime})$ transition. Because the transition probability $\mathbf{p}_{sa}$ is uniquely specified by $\bm{\theta}$, the authors simplify ambiguity sets for $\bm{\theta}$, rather than $\mathbf{p}_{sa}$, to maintain the structure of $\mathbf{p}_{sa}$. More specifically, they consider ambiguity sets with $s$-rectangular structure $\Theta=\bigotimes_{s\in\mathcal{S}}\Theta_{s}$ where $\Theta_{s}\subseteq\mathbb{R}^{k},\mathrm{~{}}\forall s\in\mathcal{S}$, and then reformulate the RMDP as:
$\max_{\pi}\min_{\bm{\theta}\in\Theta}\mathbb{E}_{\mathbf{p}sa}^{\pi}\left[\left.\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\right|\bm{\theta}\right].$ (9)
For $s\in\mathcal{S}$, $\Theta_{s}$ results from confidence regions, assuming $N$ samples from the true measure $P^{0}$ are accessible. By creating an MLE $\bm{\hat{\theta}}$ of the underlying $\bm{\theta}^{0}$, an approximate $1-\alpha$ confidence set for $\bm{\theta}_{s}$ is given by:
$\Theta_{s}=\left\{\bm{\theta}_{s}\in\mathbb{R}^{|\mathcal{A}|}\times\mathbb{R}^{k}:\sum_{a\in\mathcal{A}}\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}\right)^{T}I_{\mathbb{E}}\left(\hat{\bm{\theta}}_{sa}\right)\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}\right)\leq\chi_{k\mathcal{A},1-\alpha}^{2}\right\}.$ (10)
Taking the non-linearity of the $P$ as functions of the parameters into consideration, the authors use a discretization $\Theta_{s}^{\prime}$ of the ambiguity set $\Theta_{s}$, and reformulate the Bellman update as a linear program with $|\Theta_{s}^{\prime}|+1$ constraints.
4 Moment-Based RMDPs
While the ideas of parametric RMPDs are natural, it is challenging in practice to determine rational parameter values range. Moreover, parametric approaches impose constraints on each probability that may be too restrictive, often resulting in overly conservative or unrealistic solutions. On the other hand, it is reasonable to assume that the decision-maker can access historical observations. It is reasonable to assume that the decision-maker can access historical observations. In this case, some moment information, such as the mean and variance, may be more readily available than credible parameter value ranges, especially in data-driven paradigms.
As the seminal paper by Scarf et al. (1957) marks the inception, moment-based ambiguity sets have been extensively studied in the literature of RO and DRO for decades (Delage & Ye 2010, Wiesemann et al. 2014, Bertsimas et al. 2019, Ninh 2021), and this concept has also been expanded into into RMDPs. However, it should be noted that the literature on RMDPs employing moment-based ambiguity sets is relatively limited compared to RO and DRO. This scarcity is attributed to the nature of RMDPs, which typically model the environment as dynamics to trigger transitions, without a distinct random variable to provide clear moment information. For instance, the transition probability matrix is defined on a probability simplex in standard RMDPs, complicating the interpretation of mean and variance.
Nevertheless, if the randomness of the decision problem is exogenous, moment-based ambiguity sets can sometimes be tailored well into DRMDPs. For example, in inventory management problems, the demand is often assumed to be an independent random variable concerning initial inventory levels and decisions, and the mean and variance of demand reflect the expected value and fluctuation of uncertainty. However, in the context of DRMDPs, it seems strange to consider transition probabilities as random variables, as their meanings are challenging to decipher in a practical sense. Hence, when detached from specific issues, we suggest carefully considering whether to include moment information into the ambiguity sets of RMDPs with transition uncertainty.
Yang (2018) study the RMDPs where an estimate of the mean and covariance matrix of the disturbance $\bm{\xi}_{t}\sim\mu_{t}$ is the only available information for each stage $t$. Let $\bm{m}_{t}\in\mathbb{R}^{q}$ and $\bm{\Sigma}_{t}\in\mathbb{R}^{q\times q}$ be the estimate of the mean and covariance matrix, respectively. With the stage-wise independent assumption implicitly, the moment-based ambiguity sets are modeled over time as
$\mathcal{P}_{t}\triangleq\left\{\mu_{t}\in\mathfrak{M}(\Xi_{t})~{}\left|~{}\begin{array}[]{l}\mu_{t}(\Xi_{t})=1,\\ |\mathbb{E}_{\mu_{t}}[\bm{\xi}_{t}]-\bm{m}_{t}|\leq\bm{\theta}_{t},\\ \mathbb{E}_{\mu_{t}}[(\bm{\xi}_{t}-\bm{m}_{t})(\bm{\xi}_{t}-\bm{m}_{t})^{\top}]\preceq\gamma_{t}\bm{\Sigma}_{t}\end{array}\right\}\right.,$ (11)
where $\bm{\theta}_{t}\in\mathbb{R}^{k}$ and $\gamma_{t}\geq 1$ are given constants that depend on the confidence in the estimates $\bm{m}_{t}$ and $\bm{\Sigma}_{t}$. From a geometric perspective, the three conditions in (11) denote the following: (i) the support of $\bm{\xi}_{t}$ is $\Xi_{t}$; (ii) the mean of $\bm{\xi}_{t}$ lies in a ball of size $\bm{\theta}_{t}$; and (iii) the centered second-moment matrix of $\bm{\xi}_{t}$ lies in a positive semi-definite cone. Notice that RMDPs with ambiguity set (11) involve an infinite-dimensional mini-max optimization problem. The authors propose a dual reformulation method and prove that there is no duality gap, in virtue of the results in Lasserre (2009). Consequently, the original problem can be transformed into a semi-infinite maximal optimization program. Suppose the objective function is affine, the semi-infinite program in the dual Bellman equation can be shown to be concave with other mild assumptions.
Song et al. (2024) consider a decision-dependent epidemic control problem. Since the transition probability depends on the stochastic epidemiological processes and control manners implemented by the policy maker, the ambiguity sets are endogenous (or decision-dependent), i.e., the transition probability depends on the action $a$ and state $s$. With $(s,a)$-rectangular assumption suggestively, the authors construct ambiguity sets for each $(s,a)$ pair:
$\mathcal{P}_{sa}\triangleq\left\{\mu_{sa}\in\mathfrak{M}_{sa}(\mathcal{S})\left|\mathbf{p}_{sa}\sim\mu_{sa},~{}\bm{\theta}_{sa}^{L}\leq\mathbb{E}_{\mu_{sa}}[\mathbf{p}_{sa}]\leq\bm{\theta}_{sa}^{U}\right\},\right.$ (12)
where $\bm{\theta}_{sa}^{L}$ and $\bm{\theta}_{sa}^{U}$ are the lower bound and upper bound of the mean of transition probability. With the ambiguity sets (12), the authors provide the reformulation of Bellman equations. Then, they adopt the linear decision rule to approximate $\bm{\theta}_{sa}^{L}$ and $\bm{\theta}_{sa}^{U}$, leading to a mixed integer programming formulation to represent each Bellman equation.
Besides purely utilizing the moment information, Yu & Xu (2015) and Chen et al. (2019) both investigate the generalized-moment-based ambiguity sets, which are also called lifted ambiguity sets (Wiesemann et al. 2014). Generalized-moment-based ambiguity sets involve probability constraints on the support and expectation constraints on the generalized moment of the ambiguous distributions. Before leveraging the powerful modeling capabilities of lifted ambiguity sets, two critical conditions must be met. Firstly, the uncertain parameters across different states must be independent, namely, $s$-rectangular. Secondly, specifically for lifted ambiguity sets, the admissible state-wise ambiguity set $\mathcal{P}_{s}$ for each $s\in\mathcal{S}$ must be represented as the union of marginal distributions across all joint distributions of $(\mathbf{p}_{s},\tilde{n}_{s})$. Here $\tilde{n}_{s}\in\mathcal{N}_{s}$ is a one-dimensional auxiliary random variable that denotes a scenario. Specifically, the RMDPs with lifted ambiguity sets for each state $s\in\mathcal{S}$ can be expressed as follows:
$\mathcal{P}_{s}\triangleq\left\{\mu\in\mathfrak{M}(\mathcal{S}\times\mathcal{N}_{s})\left|\begin{array}[]{lll}&\left(\mathbf{p}_{s},\tilde{n}_{s}\right)\sim\mu\\ \vspace{1ex}&\mathbb{E}_{\mu}\left.\left[\mathbf{p}_{s}~{}\right|~{}\tilde{n}_{s}\in\mathcal{N}_{j}\right]=\theta_{j}&\forall j\in\left[J_{s}\right]\\ \vspace{1ex}&\mu\left[\mathbf{p}_{s}\in\mathcal{D}_{n}~{}|~{}\tilde{n}_{s}=n\right]=1&\forall n\in\mathcal{N}_{s}\\ &\mu\left[\tilde{n}_{s}=n\right]=\omega_{n}&\forall n\in\mathcal{N}_{s}\\ &\text{for some }\bm{\omega}\in\Delta^{|\mathcal{N}_{s}|},\left(\theta_{j},\beta_{j}\right)\in\mathcal{U}_{j}&\forall j\in\left[J_{s}\right]\end{array}\right\}\right.\end{aligned},$ (13)
where for each $j\in[J_{s}]$, the set $\mathcal{N}_{j}\subseteq\mathcal{N}_{s}$ is a subset of scenarios, and sets $\mathcal{D}_{n}\subseteq\Delta^{|\mathcal{S}|}$.
[…]
$\mathbb{E}_{\mu}\left.\left[g_{j\tilde{n}_{s}}(\mathbf{p}_{s})~{}\right|~{}\tilde{n}_{s}\in\mathcal{N}_{j}\right]\leq\beta_{j}\quad\forall j\in\left[J_{s}\right],$
where the function $g_{j\tilde{n}_{s}}$ is convex lower semi-continuous, and the authors explore and formulate $g_{j\tilde{n}_{s}}$ as the Wasserstein distance. Additionally, they show that this enhanced formulation can be effectively modeled by algebraic modeling packages (Chen et al. 2020) and solved by off-the-shelf commercial solvers.
5 Discrepancy-Based RMDPs
Besides estimating the moment information from historical observations, the decision-maker can also obtain a nominal/reference distribution to approximate the underlying probability distribution, either directly from historical data or expert experience. If data or expert experience is considerably reliable, e.g., high-quality data and recognizable patterns of uncertainty, it is reasonable to believe that the discrepancy between the nominal distribution and the true distribution is sufficiently small. More precisely, such an ambiguity set can be depicted in the following generic form:
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}\left(\Xi,\mathcal{F}\right)\left|~{}D(\hat{P},P)\leq\theta\right\}\right.,$ (14)
where $\hat{P}$ denotes the nominal probability measure, and $D(\cdot,\cdot):\mathfrak{M}\left(\Xi,\mathcal{F}\right)\times\mathfrak{M}\left(\Xi,\mathcal{F}\right)\rightarrow\mathbb{R}_{+}\cup\{\infty\}$ is a function that measures the discrepancy between two probability measures $P,~{}\hat{P}\in\mathfrak{M}\left(\Xi,\mathcal{F}\right)$. The parameter $\theta\in[0,\infty]$, also called the level of robustness, controls the size of the ambiguity set. Same as before, we abuse the notation $\mathcal{P}$ to denote the ambiguity set of corresponding $\mathbf{p}$ or $\mu$ with respect to the probability measure $P$. In line with Rahimian & Mehrotra (2019), we refer to RMDPs with ambiguity sets like (14) as discrepancy-based RMDPs, and present several prevalent ambiguity sets in this section.
5.1 Norm Ambiguity
Norm ambiguity sets, typically constructed on a probability simplex, measure the discrepancy between two distributions using various norms. While this construction appears analogous to (Satia & Lave Jr 1973), it emphasizes the difference between the entire distributions rather than individual probabilities. Generally, norm ambiguity sets can be represented as follows.
Definition 5.1 (norm ambiguity sets)
For $s\in\mathcal{S},a\in\mathcal{A}$, the $s,a$-rectangular norm ambiguity sets can be defined as
$\mathcal{P}_{sa}\triangleq\left.\left\{\mathbf{p}\in\Delta^{|\mathcal{S}|}\right|\|\hat{\mathbf{p}}_{sa}-\mathbf{p}\|\leq\theta_{sa},~{}\theta_{sa}\geq 0\right\}.$ (15)
For $s\in\mathcal{S}$, the $s$-rectangular norm ambiguity sets can be defined as
$\mathcal{P}_{s}\triangleq\left\{\mathbf{p}_{s}=(\mathbf{p}_{s1},...,\mathbf{p}_{s|\mathcal{A}|})\in\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\left|\begin{array}[]{l}\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|},\quad\forall a\in\mathcal{A},\\ \textstyle\sum\limits_{a\in\mathcal{A}}\|\hat{\mathbf{p}}_{sa}-\mathbf{p}_{sa}\|\leq\theta_{s},\theta_{s}\geq 0\end{array}\right\}\right.,$ (16)
[…]
RMDPs with norm ambiguity sets can be easily formulated as convex programs, which makes them a preferred choice for large-scale problems. Among the various norms, the $L_{1}$-norm and $L_{\infty}$-norm have garnered significant attention in the literature.
The $L_{1}$-norm (also referred to as variation distance) ambiguity sets are the most common and earliest studied norm in norm-constrained RMDPs literature (Iyengar 2005, Petrik & Subramanian 2014). The benefits of constructing such ambiguity sets are two-fold. Firstly, this construction enables the calculation of worst-case transition probabilities to be computed via linear programs, which brings computational efficiency. Secondly, the size of ambiguity sets can be easily devised using Hoeffding-style bounds with samples in the following proposition.
Proposition 5.2 ($L_{1}$-Norm Deviation Bound (Weissman et al. 2003))
Let $n_{s,a}$ be the number of transitions from state $s$ by taking action $a$, and $\delta\in(0,1]$ be the confidence level. When the deviation bound $\theta_{s,a}$ is chosen as $\theta_{s,a}=\sqrt{\frac{2}{n_{s,a}}\log\frac{|\mathcal{S}||\mathcal{A}|2^{|\mathcal{S}|}}{\delta}}$, the underlying $P_{s,a}^{0}$ is contained in the ambiguity set with probability $1-\delta$.
Ho et al. (2018) consider $L_{1}$-norm constrained RMDPs with $(s,a)$-rectangular and $s$-rectangular assumption, respectively. For $(s,a)$-rectangular assumption, the authors propose a homotopy method, assuming that the ambiguity set emerges from the intersection of the probability simplex and a $w$-weighted $L_{1}$-norm ball, where the norm is defined as $\|x\|_{1,w}=\textstyle\sum_{i=1}^{n}w_{i}|x_{i}|$. The basic idea of homotopy methods is to trace the optimal solution to an optimization problem while increasing the value of some parameters. For $s$-rectangular assumption, the authors employ the bisection method, which has been explored in $(s,a)$-rectangular case (Nilim & El Ghaoui 2005), to solve robust Bellman update efficiently. Different from the previous work, they show there is an equivalent reformulation of the robust Bellman update, and the reformulation is easily computed via the bisection method.
While the homotopy method in $(s,a)$-rectangular case in (Ho et al. 2018) accelerates the search for the optimal solution, a linear program has to be solved for each state and each step of the value or policy iteration, which is costly for a large state space. To overcome this disadvantage, Ho et al. (2021) generalize the modified robust policy iteration in Kaufman & Schaefer (2013), proposing a partial policy iteration framework for $(s,a)$-rectangular ambiguity sets.
$L_{\infty}$-Norm is another approach to construct norm ambiguity sets, which aims to restrict the maximal deviation for each transition probability mass. Although $L_{\infty}$-norm is more natural and interpretable, there exist some challenges to employing $L_{\infty}$-norm compared to $L_{1}$-norm. Similar to Ho et al. (2018, 2021), Behzadian et al. (2021) also employ homotopy and bisection methods to design quasi-linear time complexity algorithms for $L_{\infty}$-constrained $(s,a)$-rectangular and $s$-rectangular RMDPs, respectively. However, previous results and guarantees established on the sparsity properties of $L_{1}$-norm do not hold for $L_{\infty}$-norm, leading to more complex algorithms.
An interesting result for the choice between $L_{1}$-norm and $L_{\infty}$-norm is shown in Russel et al. (2019). From the perspective of value functions, they claim that the choice of set shape is dominantly driven by the structure of the value function, e.g., an $L_{\infty}$-constrained set is likely to work better than the $L_{1}$-constrained set when the value function is sparse.
5.2 $\phi$-Divergence Ambiguity
While norm ambiguity sets offer advantages in computational efficiency and modeling convenience, they are limited by their ability to provide probabilistic information and guarantees. As a result, there is growing interest in ambiguity sets designed with probabilistic metrics. $\phi$-divergence is one of the most popular class probabilistic metrics with tractable and desirably statistical properties. Here, we define general $\phi$-divergence ambiguity sets formally.
Definition 5.3 ($\phi$-divergence ambiguity)
A $\phi$-divergence ambiguity set is defined via
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}(\Xi,\mathcal{F})\left|~{}D^{\phi}(P,\hat{P})\leq\theta\right\}\right.,$ (17)
where $D^{\phi}(P,\hat{P}):=\int_{\Xi}\phi(\frac{\mathrm{d}P}{\mathrm{d}\hat{P}})\mathrm{d}\hat{P}$ is the similarity measure function whose concrete representation depends on divergence function ϕ():++{+}:italic-$\phi(\cdot):\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\cup\{+\infty\}$. The divergence function $\phi(\cdot)$ is convex, and satisfies $\phi(1)=0$, $00\phi(\frac{0}{0}):=$ 0 and $0\phi(\frac{a}{0}):=a\lim_{t\rightarrow\infty}\frac{\phi(t)}{t}$ for $a>0$ (Bayraksan & Love 2015, Pardo 2018).
Despite numerous $\phi$-divergences, only some divergences are widely accepted due to well-defined mathematical properties for robust optimization. Table 1 presents a list of typical divergence functions in probability style, where the second column $\phi(\cdot)$ is the corresponding $\phi$-divergence function, the third column $D^{\phi}(\mathbf{p}_{sa},\mathbf{q}_{sa})$ is the representation of $\phi$-divergence in probabilities style, and last column is the conjugate function of $\phi(\cdot)$.
Table 1: Typical $\phi$-divergence functions and their conjugates $\phi^{*}(a)$(Rahimian & Mehrotra 2019)
Divergence $\phi(t),t\geq 0$ $D^{\phi}(\mathbf{p}_{sa},\mathbf{q}_{sa})$ $\phi^{*}(a)$
Kullback-Leibler $t\log t-t+1$ $\sum_{s^{\prime}}p_{sas^{\prime}}\log(\frac{p_{sas^{\prime}}}{q_{sas^{\prime}}})$ $e^{a}-1$
Burg entropy $-\log t+t-1$ $\sum_{s^{\prime}}q_{sas^{\prime}}\log(\frac{q_{sas^{\prime}}}{p_{sas^{\prime}}})$ $-\log(1-a),a<1$
$J$-divergence $(t-1)\log t$ $\sum_{s^{\prime}}\left(p_{sas^{\prime}}-q_{sas^{\prime}}\right)\log(\frac{p_{sas^{\prime}}}{q_{sas^{\prime}}})$ No closed form
$\chi^{2}$-distance $(t-1)^{2}/t$ $\sum_{s^{\prime}}\frac{(p_{sas^{\prime}}-q_{sas^{\prime}})^{2}}{p_{sas^{\prime}}}$ $12-2\sqrt{1-a},a<$ 1
Modified $\chi^{2}$-distance $(t-1)^{2}$ $\sum_{s^{\prime}}\frac{(p_{sas^{\prime}}-q_{sas^{\prime}})^{2}}{q_{sas^{\prime}}}$$\begin{cases}-1&a<-2\\ a+\frac{a^{2}}{4}&a\geq-2\end{cases}$
Variation distance $|t-1|$ $\sum_{s^{\prime}}\left|p_{sas^{\prime}}-q_{sas^{\prime}}\right|$$\begin{cases}-1&a\leq-1,\\ a&-1\leq a\leq 1\end{cases}$
Hellinger distance $(\sqrt{t}-1)^{2}$ $\sum_{s^{\prime}}\left(\sqrt{p_{sas^{\prime}}}-\sqrt{q_{sas^{\prime}}}\right)^{2}$ $\frac{a}{1-a},a<1$
Kullback-Leibler (KL)-divergence is the most prevalent $\phi$-divergence in both DRO and RMDPs. As the cornerstone of modern RMDPs, Iyengar (2005) and Nilim & El Ghaoui (2005) both explore KL-divergence ambiguity sets to depict statistical uncertainty. By employing the Lagrangian duality, they reformulate the inner maximize problem in dual form,
$\min_{\lambda>0}~{}\lambda\log\left(\sum\limits_{s^{\prime}}\hat{p}_{sas^{\prime}}\exp\left(\frac{V_{t}(s^{\prime}|s,a)}{\lambda}\right)\right)+\theta\lambda,$ (18)
where $\lambda$ is the Lagrangian multiplier and $\hat{p}_{sas^{\prime}}$ is the nominal probability of transition $(s,a,s^{\prime})$. As the problem is reduced to a one-dimensional convex program, Nilim & El Ghaoui (2005) develop a bisection algorithm to obtain an $\epsilon$-optimal policy solved in polynomial time. To contribute to the marginally theoretical results, Iyengar (2005) proposes a conservative approximation ambiguity set in probability measure style,
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}(\mathcal{S})\left|\sum\limits_{s\in\mathcal{S}}\frac{(P(s)-\hat{P}(s))^{2}}{\hat{P}(s)}\leq\theta\right\}\right.,$ (19)
and show that this approximation allows us to solve the inner optimization problem more efficiently.
Caro & Das Gupta (2022) incorporate findings from Iyengar (2005) and Nilim & El Ghaoui (2005) into robust multi-armed bandit (RMAB) problem. With KL-divergence ambiguity sets, they formulate the robust Bellman equation and introduce a heuristic policy to circumvent evaluating indices, which is a well-known policy in MAB but costly for RMAB. To counteract the curse of dimensionality, they propose an approximation method. This method begins with simulating the policy based on the nominal probabilities’ point estimates and defines the states that are not encountered during the simulation as low-priority. The expected reward of low-priority states is truncated to the immediate reward. Thereby, robust value iteration is applied exclusively to the high-priority states, significantly reducing computational costs.
Similarly inspired by Iyengar (2005) and Nilim & El Ghaoui (2005), Liu et al. (2022) design a distributionally robust Q-learning algorithm with the KL divergence ambiguity sets. In virtue of the strong duality lemma from distributionally robust optimization under KL-perturbation (Hu & Hong 2013), the authors derive the distributionally robust Q-function, which transforms the infinite-dimensional problem into a finite-dimensional dual representation [^2] We display the distributionally robust Q-function without reward uncertainty and slightly modified from the original version of the paper.:
$Q(s,a):=\sup_{\alpha,\beta\geq 0}\left\{R(s,a,s^{\prime})-\alpha\theta-\gamma\beta\log\left(\mathbb{E}_{\hat{\mathbf{p}}_{s,a}}\left[\exp\left(-\frac{\max_{b\in\mathcal{A}}Q(s^{\prime},b)}{\beta}\right)\right]\right)-\gamma\beta\theta\right\}.$ (20)
The authors claim that the plug-in estimator of the distributionally robust Bellman Operator is biased because of the non-linearity. Hence, they leverage the multi-level Monte-Carlo scheme (Blanchet & Glynn 2015) to estimate the unbiased dual value using samples from the simulator and show that the proposed algorithm converges asymptotically to the optimal distributionally robust problem.
$\chi^{2}$-divergence, also known as $\chi^{2}$-test, is also commonly used to model the ambiguity sets in RMDPs. Compared to KL-divergence, $\chi^{2}$- divergence possesses several advantageous properties, including symmetry and reduced sensitivity to minor differences. Hanasusanto & Kuhn (2013) consider a data-driven stochastic control problem with continuous state and action spaces, where the transition probability matrix is estimated via Nadaraya-Watson (NW) kernel regression, as the nominal distribution. For instance, given $N$ observation histories $\{\xi^{i}_{t}\}^{T}_{t=1},i\in[N]$, the conditional expectation can be estimated as
$\mathbb{E}[V_{t+1}(s_{t+1},{\xi}_{t+1})~{}|~{}{\xi}_{t}]\approx\sum_{i=1}^{N}p_{ti}({\xi}_{t})V_{t+1}(s_{t+1}^{i},{\xi}_{t+1}^{i}),$ (21)
where $p_{ti}({\xi}_{t})=\frac{\mathrm{K_{NW}(\xi_{t}-\xi^{i}_{t})}}{\sum_{j=1}^{N}\mathrm{K_{NW}}(\xi_{t}-\xi^{j}_{t})},~{}i\in[N],t\in[T]$ and $\mathrm{K_{NW}}(\cdot)$ is the NW kernel density function. To mitigate the poor out-of-sample performance of the data-driven dynamic programming approach under sparse training data, the authors construct $\chi^{2}$-distance ambiguity sets. By general dual theorem and piece-wise linear or convex quadratic assumption of value-to-go function, they reduce the original problem to a tractable second-order cone program which can be solved efficiently with interior point algorithms. Instead of finding optimal policies, they design a robust data-driven dynamic program to approximate the value-to-go functions via interpolation.
Also utilizing $\chi^{2}$-divergence, Klabjan et al. (2013) directly use historical data to construct ambiguity sets in a goodness-of-fit test paradigm. Let the number of observations falling within the i-th bin be a function of a realized random variable $\xi_{t}\sim\mu_{t}$ as $N_{t,i}(\xi_{t})$, and the total number $n_{t}(\xi_{t})=\sum_{i}N_{t,i}(\xi_{t})$. The ambiguity set consists of the distributions $\mu_{t}$ which satisfy
$\sum_{i}\frac{(N_{t,i}(\xi_{t})-n_{t}(\xi_{t})\cdot\mu_{t}(i))^{2}}{n_{t}(\xi_{t})\mu_{t}(i)}\leq\theta,\quad t=1,...,T,$ (22)
where $\hat{\mu}_{t}(i)$ denotes the estimated probability that the number of observations falling within the $i$-th bin, and parameter $\theta$ controls how close the observed sample data is to the estimated expected number of observations according to the fitted distribution $\mu_{t}$.
Notably, the literature on RMDPs with $\phi$-divergence ambiguity sets has recently paid more attention to sample complexity and efficient algorithm design in RL. In these aspects, there are also interesting and impressive results. We briefly review several advanced literature in the following.
For sample complexity analysis, Wang, Si, Blanchet & Zhou (2023) extend Liu et al. (2022), and first provide sample complexity result for the model-free robust RL problem. Shi et al. (2023) investigate RMDPs with the uncertainty set measured via total variation distance. Under mild conditions, they prove that RMDPs in this case are more sample-efficient than standard MDPs via model-based algorithms. For efficient algorithms design, Grand-Clément & Kroer (2021) attempt to replace the original value function with the first order approximation in robust Bellman update, while Ho et al. (2022) employ bisection method to solve robust Bellman update for commonly $\phi$-divergence, e.g., KL-divergence, Burg entropy, Variation distance, and $\chi^{2}$ distance.
5.3 Wasserstein Distance Ambiguity
Although $\phi$-divergence has indeed performed well in some cases, there are some obvious limitations to obstacle the flexibility of modeling with $\phi$-divergence, e.g. only working on the same support and not satisfying the triangle inequality. Nowadays, Wasserstein distance is deemed as a good substitute for $\phi$-divergence. Wasserstein distance is a particular case of optimal transport discrepancies, which computes the minimal cost of transporting the masses between two distributions. Based on the definition of Wasserstein distance, the Wasserstein distance ambiguity is defined in the following.
Definition 5.4 (Wasserstein Distance Ambiguity)
A Wasserstein ambiguity sets can be defined via
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}(\Xi,\mathcal{F})\left|~{}\mathcal{W}(P,\hat{P})\leq\theta\right\}\right..$ (23)
Wasserstein distance $\mathcal{W}(P,\hat{P}):=\min\limits_{\Gamma\in\mathfrak{M}\times\mathfrak{M}}\left\{\int_{\Xi\times\Xi}c(\xi,\zeta)\Gamma(\mathrm{d}\xi,\mathrm{d}\zeta):~{}\Pi^{1}_{\#}\Gamma=P,\Pi^{2}_{\#}\Gamma=\hat{P}\right\}$, where $c(\xi,\zeta)$ is the cost function of moving the mass from $\xi$ to $\zeta$, $\Pi^{i}_{\#}\Gamma$ denote the $i$-th marginal distribution of $\Gamma$, and $\Gamma(\cdot,\cdot)$ is the joint distribution of $P$ and $\hat{P}$, also called transport plan.
In addition to its representation in joint distribution form, Wasserstein distance also has a dual representation, which allows for flexibility in modeling in the context of RMDPs.
Proposition 5.5
Wasserstein distance has a dual representation due to Kantorovich duality (Villani 2009), and we can represent Wasserstein distance equivalently as
$W(P,\hat{P})=\sup_{\begin{subarray}{c}\psi\in L^{1}(P)\\ \varphi\in L^{1}(\hat{P})\end{subarray}}\left\{\int_{\Xi}\psi(\xi)P(\mathrm{d}\xi)+\int_{\Xi}\varphi(\zeta)\hat{P}(\mathrm{d}\zeta):~{}\psi(\xi)+\varphi(\zeta)\leq c(\xi,\zeta)~{}\forall\xi,\zeta\in\Xi\right\},$ (24)
where $L^{1}(P)$ denotes the $L^{1}$ space of $P$-measurable functions.
Prominent results from recent Wasserstein DRO literature (Esfahani & Kuhn 2018, Blanchet & Murthy 2019, Gao & Kleywegt 2023) provide normative reformulations and finite-sample guarantees, facilitating to construct ambiguity sets in a data-driven manner and inspire the research on RMDPs with Wasserstein ambiguity sets.
Yang (2017) first investigates the Wasserstein RMDPs. The author constructs the Wasserstein ambiguity sets over each state $s\in\mathcal{S}$, falling within $s$-rectangularity implicitly, and derives corresponding distributionally robust Bellman equations by applying the dynamic programming principle,
$V_{t}(s)=\sup_{\pi\in\Delta^{|\mathcal{A}_{s}|}}\inf_{\mu\in\mathcal{P}_{s}}\int_{\Xi_{s}}\sum_{a\in\mathcal{A}_{s}}\pi(s,a)\left(R_{s}(s,a)+\textstyle\sum\limits_{s^{\prime}\in\mathbb{S}}P_{s}(s^{\prime})V_{t+1}(s^{\prime})\right)\mathrm{d}\mu(P_{s},R_{s}),$ (25)
where $(P_{s},R_{s})$ obeys to a joint distribution $\mu$. The author then derives an equivalent and tractable reformulation by leveraging the strong duality of the proposed Bellman equation.
As the extension of Yang (2017), Yang (2020) considers a continuous-state setting. Besides providing the tractable reformulation with no duality gap and iteration algorithms, the author tailors more specific properties from Wasserstein DRO literature to characterize proposed Wasserstein RMDPs. Under some mild assumptions, the optimal policy of the worst-case is deterministic and stationary, and its structure can be expressed explicitly as in Gao & Kleywegt (2023). Moreover, a probabilistic out-of-sample performance guarantee with mild assumptions is established as in Esfahani & Kuhn (2018).
Subsequently, Ramani & Ghate (2022) consider RMDPs with general distance metrics, such as total variation (TV), Burg, Hellinger, and Wasserstein distance. The authors focus on the ambiguity sets where any state-action-stage triplet is a ball, with $(s,a)$-rectangular assumption to maintain generality. They prove the optimal values of the RMDPs converge almost surely to the true optimal value as the sample size diverges to infinity, and also provide finite-sample probabilistic performance guarantee on out-of-sample values.
Besides resorting to general dynamic programming methods, there is increasing literature combining Wasserstein distance with RL algorithms. Different from previous works that merely employ Wasserstein ambiguity sets to obtain robust actions (Kandel & Moura 2020), Neufeld & Sester (2022) design a robust Q-learning algorithm to directly construct ambiguity sets, established above $(s,a)$-rectangular assumption implicitly. By introducing $\lambda c$-transform, which is proved to be equivalent to the duality of the value objective function, the authors reformulate the Q-value update as a regularized form.
Hou et al. (2020) design a Wasserstein Robust Advantage Actor-Critic algorithm (WRAAC), and leverage the strong duality proof (Blanchet & Murthy 2019) to convert the original problem into a finite-dimensional ordinary risk-aware RL problem. Through the moderated optimal Bellman equation, they prove the existence of robust optimal policies and provide the theoretical analysis of the performance of optimal policies. Based on the above theoretical foundations, the authors develop the WRAAC algorithm via two neural networks to estimate value function and approximate the action policy. WRAAC is more practical than the Wasserstein robust value iteration proposed by Yang (2017), as it can approximate nominal distribution at each step without estimating enough transition kernels.
6 Beyond Rectangularity: Coupled Uncertainty Modelling
While rectangular assumptions bring tractability, such rectangular uncertainty sets are overly conservative in modeling uncertainty and are not always appropriate in practice (Iyengar 2005). The following example demonstrates the benefits of relaxing the rectangular assumption.
Example 6.1 (Health Evolution, Goyal & Grand-Clément (2023))
A Markov model may be used to describe the health evolution of a patient. The state $s\in\mathcal{S}$ represents the health condition, and the action $a\in A$ represents the treatment. The transition probabilities $\mathbf{p}_{sa}$ represent the dynamics of the health evolution across different health conditions given a treatment. In this context, factors such as genetics and disease traits can influence how patients transition between states, resulting in uncertainty correlation in the transition probabilities across different states.
Inspired by Bertsimas & Sim (2004), Mannor et al. (2012) first propose the coupled ambiguity sets LDST motivated by the famous proverb “lightning does not strike twice”. Concretely, the ambiguity sets consist of support constraints and cardinality constraints which limit the number of times parameters can deviate from the nominal values:
$\mathcal{P}=\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\left|\mathbf{p}_{s}\in\mathcal{P}_{s},~{}\forall s\in\mathcal{S};~{}\textstyle\sum\limits_{s\in\mathcal{S}}\bm{1}(\mathbf{p}_{s}\neq\hat{\mathbf{p}}_{s})\leq K\right\}\right..$ (26)
The authors consider two models: a non-adaptive model, which can be viewed as a $T$ single-stage game; and an adaptive model, where both the decision maker and the nature are aware of the number of “deviated visits” over time and modify their strategies. Although the authors provide similar probabilistic guarantees for the two models, solving non-adaptive models is shown to be NP-hard in general and is tractable with only reward uncertainty. In contrast, with the deviation information aware, adaptive models can be tractable via augmenting the state space as
$\overline{\mathcal{S}}_{K}=\mathcal{S}\times[0:K],~{}\overline{\mathcal{S}}_{N}=\mathcal{S}\times[0:K]\times\mathcal{A},~{}\overline{\mathcal{S}}=\overline{\mathcal{S}}_{K}\bigcup\overline{\mathcal{S}}_{N},$ (27)
where $\overline{\mathcal{S}}_{K}$ denotes the states of decision maker, and $\overline{\mathcal{S}}_{N}$ denotes the states of the nature. For each $s\in\mathcal{S}$, $k\in[0:K]$, the decision-maker’s state $(s,k)$ denote the original state $s$ and deviation information $k$. Suppose the decision-maker takes action $a\in\mathcal{A}$, the next state will be $(s,k,a)$ for nature. With the augmented state space $\overline{\mathcal{S}}$ and the Nash equilibrium construction, the adaptive models are shown to be solved in polynomial time via backward induction or value iteration. Besides its tractability under non-rectangular settings, another key benefit of LDST is that it requires no distribution information.
Adhering to the same inspiration, Mannor et al. (2016) further generalizes LDST ambiguity sets to the so-called $k$-rectangular ambiguity sets, treating the LDST ambiguity set as a special case. To ensure universality, the authors introduce the definition of conditional projection which is used for $k$-rectangular ambiguity sets.
Definition 6.2 (Conditional Projection)
Let $\mathcal{S}^{\prime}\subset\mathcal{S}$ be a nonempty subset of the states, the projection of $\mathcal{P}$ to $\mathcal{S}^{\prime}$ is defined as
$\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}\triangleq\left\{P_{\mathcal{S}^{\prime}}~{}\left|~{}\exists P_{\mathcal{S}\setminus\mathcal{S}^{\prime}}\colon(P_{\mathcal{S}\setminus\mathcal{S}^{\prime}},P_{\mathcal{S}^{\prime}})\in\mathcal{P}\right\}\right..$ (28)
Given $P_{\mathcal{S}^{\prime}}\in\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}$, the projection of $\mathcal{P}$ to $\mathcal{S}\setminus\mathcal{S}^{\prime}$ conditioned on $P_{\mathcal{S}^{\prime}}$ is defined as
$\mathcal{P}_{\mathcal{S}\setminus\mathcal{S}^{\prime}}(P_{\mathcal{S}^{\prime}})\triangleq\{P_{\mathcal{S}\setminus\mathcal{S}^{\prime}}\mid(P_{\mathcal{S}\setminus\mathcal{S}^{\prime}},P_{\mathcal{S}^{\prime}})\in\mathcal{P}\}.$ (29)
Equation (29) could be interpreted as the intersection of $\mathcal{P}$ with affine subspace where $P_{\mathcal{S}^{\prime}}^{0}$ equals $P_{\mathcal{S}^{\prime}}$ from geometric perspective. Without rectangular assumption, $\mathcal{P}_{\mathcal{S}^{*}}(P_{\mathcal{S}^{\prime}})$ obviously depends on $P_{\mathcal{S}^{\prime}}$ and $\mathcal{P}^{\prime}$ for any $\mathcal{S}^{*}\subseteq\mathcal{S}$. Hence, the $k$-rectangular ambiguity sets can be defined rigorously with conditional projection..
Definition 6.3 ($k$-rectangular ambiguity sets)
For any $\mathcal{S}^{*}\subseteq\mathcal{S}$, a class of conditional projection sets $\mathfrak{P}_{\mathcal{S}^{*}}$ is defined as
$\mathfrak{P}_{\mathcal{S}^{*}}\triangleq\left\{\mathcal{P}_{\mathcal{S}^{*}}(P_{\mathcal{S}^{\prime}})\colon\forall\mathcal{S}^{\prime}\subseteq\mathcal{S}\setminus\mathcal{S}^{*},\forall P_{\mathcal{S}^{\prime}}\in\operatorname{Proj}_{\mathcal{S}^{\prime}}\mathcal{P}\right\},$ (30)
then we call it $k$-rectangular if $|\mathfrak{P}_{\mathcal{S}^{*}}|\leq k$.
The state sets $\mathcal{S}^{\prime}$ can be viewed as the historical observations, and the ambiguity sets conditioned on any $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ contains no more than $k$ sets, where each subset is corresponding to one deviation. Note that (30) is generalized from (26), both the probabilistic guarantee and tractability proof adhere to Mannor et al. (2012) in a similar way.
Different from restricting the cardinality of deviations, Wiesemann et al. (2013) resort to “linear decision rules”, an approach adopted in the RO literature (Ben-Tal et al. 2009), using constant or (piece-wise) linear functions to approximate the value-to-go function. For a non-rectangular ambiguity set $\mathcal{P}$, the authors first construct $\mathcal{\bar{P}}:=\bigotimes_{s\in\mathcal{S}}\mathcal{P}_{s}$ as the smallest $s$-rectangular ambiguity set that contains $\mathcal{P}$. Then, they propose an algorithm to optimize the linear approximations of the value-to-go function over the rectangularized ambiguity set and yield an optimal but overly conservative solution in polynomial time.
Motivated by Example 6.1, Goh et al. (2018) and Goyal & Grand-Clément (2023) emphasize the transition probability is always affected by environment features. Thanks to big data technology, the authors can develop a new framework to explicitly characterize the relationship between features/factors and transition probability, and they call the modeling framework a factor uncertainty model. A factor matrix ambiguity set is defined as
$\left.\mathbf{p}=\left\{\left(\textstyle\sum\limits_{i=1}^{r}u_{sa}^{i}w_{i,s^{\prime}}\right)_{sas^{\prime}}\right|\bm{W}=(\bm{w}_{1},\ldots,\bm{w}_{r})\in\mathcal{W}\subseteq\mathbb{R}^{S\times r}\right\},$ (31)
where coefficients $u_{1},...,u_{S}$ are fixed and known in $\mathbb{R}_{+}^{r\times\mathcal{A}}$ and underlying factors $\mathcal{W}\in\mathbb{R}_{+}^{\mathcal{S}\times r}$ is a convex, compact subset such that
$\sum\limits_{i=1}^{r}u_{sa}^{i}=1,\forall(s,a)\in\mathcal{S}\times\mathcal{A},~{}\textstyle\sum\limits_{s^{\prime}=1}^{S}w_{i,s^{\prime}}=1,\forall i\in[r],$ (32)
We would like to highlight that, akin to the approach outlined in Wiesemann et al. (2013), this model represents another instance of the utilization of linear decision rules. However, in this context, linear decision rules are applied to approximate the transition probability.
Tirinzoni et al. (2018) provide a distinct scope as a robust improvement. The authors consider a new policy $\pi$, which can achieve robust improvement compared to base policy $\tilde{\pi}$. Being aware that state-action pairs are inefficient in characterizing the underground truth under a non-rectangular setting, they consider incorporating some crucial triple $(s,a,s^{\prime})$ into ambiguity, which can be thought moment information about these important transitions. Concretely, the authors define a feature function $\phi(\cdot):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$, inspired from inverse reinforcement learning literature, which is always chosen as an indicator function. And then, they construct ambiguity sets as
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\left|~{}\mathbb{E}_{\mathbf{p}}^{\tilde{\pi}}\left[\textstyle\sum\limits_{t=1}^{T}\phi(s_{t},a_{t},s_{t+1})\right]=\frac{1}{N}\textstyle\sum\limits_{i=1}^{N}\textstyle\sum\limits_{t=1}^{T}\phi(s_{t}^{(i)},a_{t}^{(i)},s_{t+1}^{(i)})\right\}\right..$ (33)
The change from $(s,a)$ pair to $(s,a,s^{\prime})$ triple essentially involves the comprehensive incorporation of entire trajectories.
Recently, Wang, Ho & Petrik (2023) propose a double-loop algorithm for RMDPs, where the outer-loop convergence does not rely on any rectangular assumption. However, in their framework, it is still difficult to solve without any rectangular assumptions for the inner loop.
7 Other Related Literature and New fashion RMDPs
What we have introduced so far still falls within the traditional min-max (max-min) framework, emphasizing the pursuit of an optimal solution to hedge against the (conditional) worst-case. However, the worst-case performance of these approaches is sometimes overly conservative, offering little practical insight into the actual performance of reliable policies. In the RMDPs community, new fashionable modeling approaches have emerged, most inspired by recent advancements in RO and DRO. We will introduce these emerging frameworks and related literature as an extension for future research directions.
7.1 Satisfactory Regime
RMDPs can mitigate the impact of perturbations, however, the optimized worst-case performances are often too pessimistic and not easy to offer insights and interpretation for decision-makers (Long et al. 2023). To address this issue, the satisfactory RMDPs have emerged. In the satisfactory regime, there are two mainstream formulations: Safe Policy Improvement (SPI) and Robust Satisfying (RS).
SPI is established in Batch Reinforcement Learning (Batch RL), where only limited trajectories and a behavioral/baseline policy $\pi_{B}$ to collect the trajectories are accessible. The goal of SPI is obtaining a policy that is guaranteed to perform at least as well as $\pi_{B}$. The general SPI problem can be described as follows.
Definition 7.1 (Safe Policy Improvement Problem)
Given the uncertainty set $\mathcal{P}$ and a baseline policy $\pi_{B}$, find a maximal $\theta>0$ such that there exists a policy $\pi$ for which $\rho(\pi,\mathbf{p})\geq\rho(\pi_{B},\mathbf{p})+\theta,~{}\forall\mathbf{p}\in\mathcal{P}$, which is equivalent to the following optimization problem:
$\pi_{S}\in\arg\max_{\pi}\min_{\mathbf{p}\in\mathcal{P}}\left[\rho(\pi,\mathbf{p})-\rho(\pi_{B},\mathbf{p})\right].$ (34)
Seminar paper by Ghavamzadeh et al. (2016) considers $L_{1}$-norm uncertainty sets where the discrepancy is bounded by an error function $e(s,a)$. The authors prove that problem (34) may be purely randomized and is an NP-hard problem, even though the uncertainty set is $(s,a)$-rectangular. Due to the intractable formulation, the authors propose a simple and practical approximate algorithm in which either improvement is significant or deviation (error function) is small; otherwise, the baseline policy $\pi_{B}$ is recommended in all other cases.
Subsequently, Laroche et al. (2019) and Simão et al. (2020) coincidentally generalize SPI as SPI with baseline bootstrapping to mitigate the unreliability due to the large state space and restrictive assumptions. The core procedure is to limit state-action pairs into a bootstrapped set $\mathcal{B}$, instead of searching for the analytic optimum over the whole space. In virtue of the statistical properties of bootstrapping, Laroche et al. (2019) proposed a computationally efficient algorithm that provides a PAC-style guarantee of policy improvement with high confidence, and Simão et al. (2020) provide SPI guarantees over the true baseline even without direct access to it.
RS is a novel framework developed in RO and DRO (Long et al. 2023) that transforms the traditional optimization problem of seeking the optimum into attaining an acceptable target specified ahead. Due to the outstanding interpretability and complete theory of RS, Ruan et al. (2023) extend it into the MDPs version in recent. As opposed to traditional RMDPs with hard constraints to limit the distinction among distributions, RS MDPs impose soft constraints for all other transition distributions. Leveraging the theorems from Long et al. (2023), the authors formulate the original problems as a conic program.
The decision variables reflect the magnitude of constraint violation incurred by the distance between the ambiguous transition kernel and the empirical one. Besides, they transform the original problems into a min-max form and propose a primal-dual algorithm to solve the min-max problem efficiently.
7.2 Online and Bayesian Regime
Another significant drawback of traditional RMDPs lies in their inability to “learn” from uncertainty, which easily causes overly conservative policies due to the extensive uncertainty sets. To tackle this challenge and infuse the adaptive learning characteristic of RL, there is a growing trend to incorporate online learning or Bayesian concepts into the design of uncertainty/ambiguity sets, Bellman updates, environmental configurations, and other facets of RMDPs. Such efforts are expanding the scope of RMDPs and injecting fresh dynamism into the field.
Out of consideration for limiting the size of uncertainty set, Lim et al. (2013) assume there is a unique but unknown subset $\mathcal{Z}$ of state–action pairs behaving adversarially while all others are stochastic, as in non-robust MDPs. In essence, they propose a two-step framework: divide pairs into two categories, then compute respectively. Motivated by Bubeck & Slivkins (2012) that investigate a similar scenario in a multi-armed bandit setting, the authors implement a statistical hypothesis test, termed consistency check, after executing $(s,a)$ every time. All state-action pairs that pass the consistency check are stochastic, while the pairs that do not pass the check will consist of $\mathcal{Z}$. Combining the consistency check with UCRL2, a well-known online algorithm that treats every state-action pair as stochastic (Auer et al. 2009), the authors develop two algorithms for finite-horizon and infinite-horizon cases, respectively. Adhering to the optimistic property of UCRL2, the results of the proposed algorithms are less conservative than the standard RMDPs.
Sharing similar concerns but adopting distinct methodologies, Petrik & Russel (2019) observe that reliance on concentration inequalities results in conservative ambiguity sets, and then the authors introduce Bayesian inference to construct tight ambiguity sets. In the paper, the authors focus on a safe return measured by Value-at-Risk (VaR) and aim to improve the lower bound when the value function is unknown. Under Bayesian assumptions, the transition probabilities $\mathbf{p}_{sa}$ is a random variable with a prior distribution. Specifically, they propose the robustification with sensible value functions (RSVF) concept, in which the ambiguity set $\mathcal{P}_{sa}$ is derived from a set of possible value functions $\mathcal{V}$. Meanwhile, given any $v$, they approximate the optimal ambiguity set as
$\mathcal{K}_{sa}(v)=\left\{\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}\left|\mathbf{p}_{sa}^{\top}v\leq\operatorname{VaR}_{\mathbf{p}^{0}}^{\theta}[\mathbf{p}_{sa}^{0\top}v]\right\}\right.,$ (35)
where $\mathbf{p}^{0}$ serves as the nominal probability.
As augmenting value function into $\mathcal{V}$, the algorithm updates $\mathcal{P}_{sa}$ and currently optimal value function $\hat{v}$. If the intersection $\mathcal{K}_{sa}(\hat{v})\cap\mathcal{P}_{sa}\neq\emptyset$ for all $(s,a)$ pairs, the algorithm deems the current policy and value to be safe. For theoretical contributions, they provide a finite iteration probabilistic guarantee that is beneficial from their special formulation and empirically compare the safe estimates from RSVF and other non-Bayesian methods. Empirical results demonstrate that RSVF indeed provides much tighter ambiguity sets and safe returns.
Unlike the aforementioned works, Derman et al. (2020) offer a posterior perspective to design Bellman update for Bayesian RMDPs. Collecting an observation history $\mathcal{T}:=\{(s_{1},a_{1}),...,(s_{t},a_{t})\}$ induced by policy $\pi$ and a predetermined confidence level $\theta_{sa}>0$ for each pair, the posterior uncertainty sets are constructed over time:
$\widehat{\mathcal{P}}^{t}_{sa}\triangleq\left\{\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}\big{|}\|\mathbf{p}_{sa}-\bar{\mathbf{p}}_{sa}\|_{1}\leq\theta_{sa}\right\},$ (36)
where $\bar{\mathbf{p}}_{sa}:=\mathbb{E}[\mathbf{p}_{sa}|\mathcal{T}]$ acts as the nominal transition probability, and the overall ambiguity set satisfies $(s,a)$-rectangular assumption. Once the posterior uncertainty set is established, a posterior over robust Q-values at stage $t$ can then be computed as:
$\widehat{Q}^{t}(s,a)=R(s,a)+\gamma\inf_{\mathbf{p}\in\widehat{\mathcal{P}}^{t}_{sa}}\textstyle\sum\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\pi^{t}(a^{\prime}\mid s^{\prime})p_{sas^{\prime}}\widehat{Q}^{t+1}(s^{\prime},a^{\prime}).$ (37)
Equation (37) is termed as Uncertainty Robust Bellman Equation (URBE). This uncertainty set is constructed for each episode based on the observed data from all previous ones.
Slightly modifying the procedures in O’Donoghue et al. (2018), the authors compute posterior variance $\omega_{sa}^{t}$ of $\widehat{Q}^{t}(s,a)$ in (37), and approximate the posterior over robust Q-values as normal distribution $\mathcal{N}(\bar{Q},\mathrm{diag}(\bm{\omega}))$, where $\mathrm{diag}(\bm{\omega})$ is the solution of URBE and $\bar{Q}$ the conditional expectation of $\hat{Q}$. This approach offers a trade-off between robustness and conservatism for robust policies. Besides, the authors propose a DQN-URBE algorithm, and they show that it is significantly faster to change dynamics online compared to existing robust techniques with fixed uncertainty sets.
Contrary to most studies that concentrate on developing innovative robust algorithms to learn from the original environment, Wang, Gadot, Kumar, Levy & Mannor (2023) introduce an RL framework designed to approximate an adversarial environment and generate adversarial data for training purposes. This approach is inspired by Kumar et al. (2023) which reveals that the adversarial distribution fundamentally alters the next-state transition probability from the nominal one. Compared to applying intricate iterations of robust methods, directly employing non-robust methods in adversarial environments will be more computationally efficient, thus addressing scalability challenges in RMDPs with high-dimensional domains. Different from the $L_{p}$-norm ambiguity sets used by Kumar et al. (2023), the authors construct KL-divergence ambiguity sets under $(s,a)$-rectangular assumption, which circumvents the issue of different supports between nominal and adversarial distribution. They also derive a closed-form expression for the adversarial distribution, depicted as a re-weighted nominal distribution. Consequently, the framework can be straightforwardly implemented by simulating the next state under the nominal dynamic and choosing it with corresponding adversarial weight. After realizing the next state, $(s,a,s^{\prime})$ tuple is added to the data buffer, and the policy is trained with data from the buffer via any non-robust RL method.
7.3 Risk, Regularization, and Robustness
Our pursuit of robustness aims to enhance the generalization performance and mitigate the uncertainty when the underlying truth is ambiguous. This concept resonates across various fields, e.g., finance and machine learning (ML), where both risk-aware and regularized formulations have seen considerable success. Notably, a body of literature attempts to bridge these concepts with RO and DRO (Shapiro 2017, Namkoong & Duchi 2017, Shapiro 2021, Shafieezadeh-Abadeh et al. 2019, Blanchet & Murthy 2019), and also RMDPs. These studies offer diverse interpretations of robustness, revealing connections that shed light on the models of Markov Decision Processes (MDPs). This understanding helps to partially alleviate the conservatism inherent in robust policies.
In terms of risk-aware MDPs, Osogami (2012) explores the equivalence between RMDPs and risk-aware MDPs with an expected exponential utility objective. By leveraging the properties of this objective, the author reformulates risk-aware MDPs into a general RMDP that aims to minimize the expected cumulative cost adjusted by the KL-divergence penalty. This penalty quantifies the discrepancy between the nominal and the worst-case distribution. Building on this foundation, Bäuerle & Glauner (2022) broaden these equivalence findings to encompass all risk-aware MDPs that employ spectral risk measures, which are known as a class of coherent risk measures like the Expected Shortfall. However, the risk measures examined are confined to the Markov risk measure (Ruszczyński 2010), known to be potentially non-gradient-dominant (Huang et al. 2021).
More recently, Zhang et al. (2023) establish the equivalence between risk-aware MDPs and a class of regularized RMDPs, including the standard RMDPs. The authors introduce an innovative formulation for risk-aware MDPs that incorporates general convex risk measures, requiring only that these measures adhere to principles of monotonicity, translation invariance, and convexity. This formulation facilitates a broader equivalence and supports the use of policy gradient methods for algorithm design, ensuring global convergence.
Similarly, the equivalence between regularization and robustness in RMDPs has been investigated recently. Derman & Mannor (2020) consider a reward-maximizing RMDP within the Wasserstein distance ambiguity set. Employing strong duality and necessary conditions, they construct a regularized value function that serves as a lower bound to the distributionally robust value. Moreover, they extend the finite-sample guarantee results of Yang (2020) to regularized MDPs, though they stop analyzing the tightness and asymptotic consistency of their approach as the sample size increases.
As a supplement and improvement of Derman & Mannor (2020), Derman et al. (2021) formally demonstrate the equivalence between regularized MDPs and RMDPs with arbitrary ball-constrained ambiguity sets, extending beyond the Wasserstein metric. They propose a comprehensive RMDP model that accounts for uncertainties in both rewards and transitions, introducing a novel extension to regularized MDPs that encompasses both policy and value regularization. Utilizing contraction and monotonicity assumptions, they successfully apply Banach’s fixed point theorem to this extended model and show that equivalence is maintained. As a byproduct, they propose efficient Bellman updates for a modified policy iteration tailored to the structure of the regularized problem. While their method solves RMDPs similarly to classical non-robust MDPs, the obtained policies can be overly conservative due to the lack of restrictions on worst-case transition probabilities to the probability simplex.
Because of theoretical equivalence and computational efficiency, the integration of risk measures and regularization into MDPs has recently gained significant momentum. Value-at-Risk (VaR) and Conditional Value-at-Risk (CVaR), as the most widely recognized risk metrics, have spurred a considerable amount of research in quantile-based MDPs. In the model-based domain, Li et al. (2022) investigate quantile MDP(QMDP) which is defined as
$\max_{\pi\in\Pi}Q_{\tau}^{\pi}\left[\sum_{t=0}^{T-1}r_{t}(s_{t},a_{t},\xi_{t})\right],$ (38)
where $Q_{\tau}(X):=\inf\{x\mid P(X\leq x)\geq\tau\}$ and $\tau\in(0,1)$ is the quantile value. The authors reformulate (38) as a max-min formulation and have pioneered the verification that the quantile-based value-to-go function is tantamount to solving a specific optimization problem for a given action. This breakthrough facilitates the application of dynamic programming techniques. Additionally, they craft efficient algorithms and conduct a detailed complexity analysis. In terms of model-free approaches, Yu & Shen (2022) adapt CVaR as a special one-step conditional risk measure to preserve time consistency and subsequently reformulate the risk-averse MDP as a risk-neutral counterpart with augmented action space and the adjusted immediate rewards. Furthermore, the authors show that the Bellman operator in this formulation is a contraction mapping, and consequently, the authors extend Q-learning results into this case and develop a risk-averse deep Q-learning algorithm.
The concept of weaving regularization into dynamic programming also has been thoroughly investigated within the reinforcement learning field. (Kaufman & Schaefer 2013, Neu et al. 2017, Kostrikov et al. 2021, Ho et al. 2021, Kumar et al. 2022). While these regularized variants of value iteration and policy iteration have been developed, Grand-Clément & Petrik (2022) first propose the convex optimization formulation of RMDPs. This innovative approach integrates an entropic regularization of the robust Bellman operator with a change of variables involving exponential and logarithmic functions. By conducting this convex reformulation, arbitrary algorithms designed for solving convex optimization problems can be incorporated with the reformulated RMDPs. This fresh perspective has paved the way for the development of innovative algorithms tailored to RMDPs.
8 Conclusion
This survey provides an overview of recent methodologies for modeling RMDPs with transition uncertainty models. We start with a framework that encompasses three primary means of utilizing information to construct ambiguity sets in RMDPs. The parametric approach is the simplest yet most restrictive way to characterize the uncertainty, usually requiring constrained support or distribution family. Moment-based ambiguity sets are more suitable to exogenous randomness and concrete problem background. As the most popular approach, discrepancy-based RMDPs match in a data-driven manner. All three mainstream discrepancy measures–norm, $\phi$-divergence, and Wasserstein distance, can determine the size of ambiguity sets that contain underlying probability measures or distributions with high probability. It is noteworthy that norm and $\phi$-divergence may be more favored in fast algorithm research, while Wasserstein distance displays more powerful performance in modeling and practice.
We also review recent efforts that depart from rectangular assumptions. Rectangularity is the key assumption in most RMDPs research, which allows decomposability to ensure that RMDPs can be solved in polynomial time. We summarize three class definitions of rectangularity and prove that solving non-rectangular RMDPs is NP-hard. There are some attempts to relax traditional rectangular assumptions. While some attempts have been made to relax the traditional rectangular assumption, these coupled uncertainty modeling approaches often compromise the flexibility of the proposed model. Besides the minimax framework, we explore related paradigms in the literature that seek to achieve similar performance to traditional RMDPs. These new paradigms in RMDPs are often represented as combinations of RL and other techniques. These works establish a bridge between RMDPs and other areas, bringing new perspectives for hedging uncertainty.